OpenStudy (anonymous):

Evaluate 9 added to the g(x) when x = 8. g(8)+9=?

OpenStudy (anonymous):

@dan815 @SamsungFanBoy

OpenStudy (dan815):

You need to look at the other information given

OpenStudy (anonymous):

okay but i dont understand this

OpenStudy (anonymous):

Ok so g=9 and x=8

OpenStudy (anonymous):

whats 9x8

OpenStudy (anonymous):

63?

OpenStudy (anonymous):

no 72

OpenStudy (anonymous):

okay ya sorry fuked up my math

OpenStudy (anonymous):

ok now we know g=72 whats 72x8?

OpenStudy (anonymous):

576

OpenStudy (anonymous):

yes Then add 9 whats 9+572

OpenStudy (anonymous):

576

OpenStudy (anonymous):

585

OpenStudy (anonymous):

thats your anser :D

OpenStudy (anonymous):

so g(8)+9=585

OpenStudy (anonymous):

Yes :D

OpenStudy (anonymous):

can you help me on 2 more :D

OpenStudy (anonymous):

what grad r u in tho

OpenStudy (anonymous):

and yes

OpenStudy (anonymous):

11/12th and thanks:)

OpenStudy (anonymous):

o pellet im in 7 lol

OpenStudy (anonymous):

omg lol it canged it to pellet

OpenStudy (anonymous):

lol yea i hate that and looks like your smarter then me lmao

OpenStudy (anonymous):

its ok its hard at times dont qwout me on that

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

i will be back in ten ok

OpenStudy (anonymous):

Evaluate 5 times the g(x) when x = 3. 5[g(3)]

OpenStudy (anonymous):

ok srry idk on this won

OpenStudy (anonymous):

lol its okay how bout this one Evaluate the product of f(x) and g(x) when x = 4. f(4)*g(4)=?

OpenStudy (dan815):

again, you are not giving us all the information

OpenStudy (anonymous):

hey dan can u check my work plz

OpenStudy (dan815):

Go back to the top of your question, and see if f(x) and g(x), they are some functions of x and should be defined

OpenStudy (dan815):

okay sure :) @The-Bruh-Frog

OpenStudy (anonymous):

thts all i know... the question exact was 29) Evaluate the product of f(x) and g(x) when x = 4. f(4)*g(4)=? \[f(x)=2x+5 , g(x)=x^2+3\]

OpenStudy (dan815):

Okay we need those functions down there, now sub in 4 into both of them

OpenStudy (dan815):

\[f(4)*g(4)=(2(4)+5)*((4)^2+3)\]

OpenStudy (anonymous):

247

OpenStudy (dan815):

You understand how to do these questions now?

OpenStudy (dan815):

can you do this question now? lets say g(x) is that function you gave up there x=8 then g(x)+9=g(8)+9 =?

OpenStudy (dan815):

you can ask a question if u are still confused, i wont bite! xD

OpenStudy (anonymous):

no its all good:) i might need your help if my teacher doesn't except what i had given him so if you would keep an eye out on if i tag you in anything:)